OpenStudy (anonymous):

simplify 2 over 5 divided by 3 and a half?

OpenStudy (anonymous):

4/35

OpenStudy (anonymous):

how did you get that show all steps

OpenStudy (anonymous):

\[\frac{\frac{2}{5}}{\frac{3}{2}} = \frac{2}{5} * \frac{2}{3} = \frac{4}{15}\]

OpenStudy (anonymous):

2/5 divided by 3 1/2

OpenStudy (anonymous):

Use LaTeX pls :)